OpenStudy (3psilon):

Trig help The following information refers to triangle ABC. Find angle B A = 42° 30', C = 120° 30', a = 4.53 ft

OpenStudy (3psilon):

What does the 30' part mean?

OpenStudy (jdoe0001):

check your trig sheet for the "law of sines"

OpenStudy (jdoe0001):

ohh, the 30' means 'minutes', there are 60minutes to a degree, and 60'' seconds to a minute, 30minutes off a total of 60 is 0.5 degree

OpenStudy (jdoe0001):

so, \(42^o30'\) can also be written as \(41^o89'60''\) :)

OpenStudy (3psilon):

Thanks man! But what do I do with the minutes

OpenStudy (3psilon):

When finding the angle

OpenStudy (jdoe0001):

make them decimals :), to work with the Law of Sines

OpenStudy (jdoe0001):

30' off a total of 60 is 0.5

OpenStudy (3psilon):

Like half of 42 ? or 42.5 degrees?

OpenStudy (jdoe0001):

how many degrees do all internal angles in a triangle sum up to?

OpenStudy (3psilon):

180

OpenStudy (3psilon):

it says 18 isn't the angle though

OpenStudy (jdoe0001):

right, you have \(\angle A\ and\ \angle C\), so A+B+C=180, so B=180-(A+C) :)

OpenStudy (jdoe0001):

tis a tiny angle by the way

OpenStudy (3psilon):

yeah I got 18

OpenStudy (3psilon):

But it wasnt right

OpenStudy (jdoe0001):

well, \(42^o30'+120^o30'\) is not the same as \(42^o+120^o\)

OpenStudy (jdoe0001):

come to think, you don't even need the law of sines :|, since all you're after is \(\angle B\)

OpenStudy (3psilon):

If it isnt the same what is it then

OpenStudy (jdoe0001):

what would you get from \(42^o30'\) making the 30' a decimal?

OpenStudy (3psilon):

42.5

OpenStudy (jdoe0001):

yeah, so B=180-(42.5+120.5)

OpenStudy (3psilon):

Thanks I got it! :_)

OpenStudy (3psilon):

:o)

OpenStudy (jdoe0001):

yw